OpenStudy (anonymous):

6(z-2)+4=10 I will give out medals to the person with the best answer! c:

OpenStudy (tkhunny):

Let's first work on your Distributive Property. Expand and remove those parentheses. Go!

OpenStudy (anonymous):

6z-12+4=10? Is that right? @tkhunny

OpenStudy (tkhunny):

Excellent. Now, combine like terms on the left-hand side.

OpenStudy (anonymous):

6z-16=10? @tkhunny

OpenStudy (tkhunny):

No good. -12 + 4 = -8 6z - 8 = 10

OpenStudy (anonymous):

Okay, but I thought combining like terms would be 12+4 and that would be 16. What am I missing here then that I don't understand?

OpenStudy (tkhunny):

One must consider the sign at all times. If you get 16, where did the "-" go? It seems magically to have vanished. If it were just -12 + 4, would you still get -16?

OpenStudy (anonymous):

Oh, I see where you are coming from now, but -12+4=-8, so no.

OpenStudy (anonymous):

So now all I got do is this: 6z-8=10 +8 +8 6z 18 6z/6=6 18/6=3 So that means that the answer is 3?

OpenStudy (tkhunny):

Good. -12 + 4 = -8 SO... 6z - 12 + 4 = 6z - 8 Looks good. Check it out in the ORIGINAL equation.

OpenStudy (anonymous):

Alright, it works. Thank you for your services and here is your medal like I promised! :)